# CL-RAG: Bridging the Gap in Retrieval-Augmented Generation with Curriculum Learning

### v1

@@ Abstract
Retrieval-Augmented Generation (RAG) is an effective method to enhance the capabilities of large language models (LLMs). Existing methods focus on optimizing the retriever or generator in the RAG system by directly utilizing the $top$-$k$ retrieved documents. However, the documents effectiveness are various significantly across user queries, i.e. some documents provide valuable knowledge while others totally lack critical information. It hinders the retriever and generator’s adaptation during training. Inspired by human cognitive learning, curriculum learning trains models using samples progressing from easy to difficult, thus enhancing their generalization ability, and we integrate this effective paradigm to the training of the RAG system. In this paper, we propose a multi-stage Curriculum Learning based RAG system training framework, named CL-RAG. We first construct training data with multiple difficulty levels for the retriever and generator separately through sample evolution. Then, we train the model in stages based on the curriculum learning approach, thereby optimizing the overall performance and generalization of the RAG system more effectively. Our CL-RAG framework demonstrates consistent effectiveness across four open-domain QA datasets, achieving performance gains of 2% to 4% over multiple advanced methods.
CL-RAG: Bridging the Gap in Retrieval-Augmented Generation with Curriculum Learning
Shaohan Wang, Licheng Zhang , Zheren Fu, Zhendong Mao University of Science and Technology of China, Hefei, China {wsh2000, zlczlc, fzr}@mail.ustc.edu.cn     zdmao@ustc.edu.cn
1 Introduction
Figure 1: Comparison between our method and previous studies, where green represents documents that are conducive to model responses, while red indicates documents that are useless or even detrimental to model responses. Figure 1(a) illustrates the previous method, which randomly samples documents of varying quality without considering the difficulty order of the documents. Figure 1(b), in contrast, presents our approach, which focuses on reordering and combining documents based on their difficulty levels for stage-by-stage training.
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities in a wide range of Natural Language Processing (NLP) tasks (Brown et al., 2020; Anil et al., 2023; Dubey et al., 2024), but they still constrained by the limitations of the knowledge embedded within their internal parameters (Roberts et al., 2020; Kandpal et al., 2023; Gao et al., 2023). Retrieval-Augmented Generation (RAG) addresses this limitation by leveraging additional knowledge retrieved from external knowledge bases. By retrieving relevant documents and incorporating them as contextual input, RAG has significantly enhanced the capabilities of existing large models in tasks such as Open-Domain Question Answering (Izacard et al., 2023; Shi et al., 2023; Yoran et al., 2023; Lin et al., 2023; Fang et al., 2024) and Natural Language Modeling (Borgeaud et al., 2022; Ram et al., 2023; Zhang et al., 2024). The overall performance of the RAG system depends on the quality of the retrieved documents and the LLMs’ ability to utilize these documents. Therefore, how to retrieve better documents and how to better utilize the retrieved documents are of vital importance in RAG research.
Existing efforts to enhance RAG systems have focused on: (1) Utilizing LLMs to evaluate document quality and guide retriever training (Shi et al., 2023; Zhang et al., 2024); (2) Training Retrieval-Augmented Language Models (RALMs) that can better utilize documents (Izacard and Grave, 2020; Yoran et al., 2023; Fang et al., 2024), or a combination of both (Lin et al., 2023) to enhance the overall performance of the RAG system. However, given a specific query, these researches directly use the $top$-$k$ retrieved documents as training samples and randomly input them into the model for training, as shown in Figure 1(a). This approach neglects the substantial quality variations among retrieved documents, where the $top$-$k$ documents for different queries may vary greatly in quality: some provide valuable knowledge while others lack critical information or even contain misleading content. These discrepancies pose varying learning challenges for the retriever and generator during training.
Inspired by cognitive science studies (Elman, 1993; Rohde and Plaut, 1999) that humans can benefit from a easy-to-difficult learning sequence, the curriculum learning (Bengio et al., 2009) suggests that deep learning models trained by the gradual introduction of more challenging samples in subsequent stages. Previous studies (Guo et al., 2018; Hacohen and Weinshall, 2019; Xu et al., 2020) have demonstrated that this easy-to-difficult strategy can effectively enhance the generalization ability of models. Therefore, we propose that the training of the RAG system should also adhere to this paradigm, which has been neglected in previous works. To bridge this gap, we propose a sample evolution strategy to construct document data with different levels of difficulty and further train the retriever and generator based on the idea of curriculum learning, as demonstrated in Figure 1(b).
In this work, we propose a Curriculum Learning based training framework for the RAG system, named CL-RAG. To the best of our knowledge, we are the first to integrate the idea of human imitation learning with RAG training, which can effectively enhance the generalization and stability of the RAG system. Specifically, we first focus on the training of the generator, constructing multiple difficulty levels of documents by rewriting query-enhanced golden documents or counterfactual distractor documents. Then we finetune a RALM in a stage-by-stage manner according to the order of difficulty. For the training of the retriever, we propose to use the well-trained RALM to assess the quality of documents and rerank them. We then construct document data from easy to difficult by gradually reducing the ranking gap between sampled documents, and propose a hierarchical training strategy to finetune the retriever.
@@
As a result, CL-RAG bring two advantages to RAG system: (1) For the generator, CL-RAG systematically transition it from merely extracting information to effectively countering potential distracting noise within documents. (2) For the retriever, CL-RAG train it in stages to progressively distinguish documents with obvious differences and then distinguish those with only slight differences. This progressive training framework has led to an overall performance and generalization improvement of the RAG system. Our contributions can be concluded as follows:
@@
  • We propose the CL-RAG training framework for the RAG system based on the concept of curriculum learning. To the best of our knowledge, this is one of the first times that CL strategy has been applied to optimize RAG systems.
@@
  • We defined the document difficulty levels for the retriever and generator based on the concept of curriculum learning and designed a complete method for constructing enhanced documents.
@@
  • We evaluated our CL-RAG framework on four popular datasets, demonstrating consistent performance gains of 2% to 4% over multiple advanced methods, thereby highlighting the superiority of our approach.
Figure 2: The overview of our CL-RAG training framework, which has into two continuous phases: (1) Training Generator: We construct multiple difficulty levels of documents and then finetune a RALM in a stage-by-stage manner. (2) Training Retriever: We use the well-trained RALM to assess documents and rerank them. We then construct document data from easy to difficult, and finetune the retriever in a stage-by-stage manner.
2 Related Work
2.1 Retrieval-augmented Generation
@@
Using documents retrieved from extended knowledge bases to enhance the capabilities of large language models (LLMs) has been proven effective in NLP tasks, including language modeling (Borgeaud et al., 2022; Ram et al., 2023; Zhang et al., 2024) and question answering (Izacard et al., 2023; Shi et al., 2023; Yoran et al., 2023; Lin et al., 2023; Fang et al., 2024). Specifically, a Retrieval-Augmented Generator (RAG) system takes a query as input and uses a retriever to retrieve relevant documents from an external knowledge base. Then, it combines the documents with the query and feeds them into the LLM to make up for the LLM’s own lack of knowledge.
Optimization of the RAG system focuses on two main areas: improving the retriever and enhancing the generator (LLM) to the RALM. Replug (Shi et al., 2023) uses KL divergence to align retriever results with LLM preferences. LLM-Embedder (Zhang et al., 2024) employs a distillation objective based on LLM rankings. These methods train the retriever to better match LLM preferences. For generator, FiD (Izacard and Grave, 2020) finetunes LLM to handle retrieved documents and queries, addressing irrelevant information. Other studies introduce noise to improve LLM robustness (Yoran et al., 2023; Fang et al., 2024). Combining the strengths of both approaches, RA-DIT (Lin et al., 2023) uses modular training to optimize the retriever and LLM separately, enhancing overall RAG system performance.
2.2 Curriculum Learning
Curriculum Learning (CL) is a machine learning strategy that mimics human learning by training models gradually, starting with simpler tasks and progressing to more complex ones. It aims to improve the model’s generalization ability and accelerate its convergence speed. Early studies have extensively investigated CL in computer vision domain and have demonstrated its advantages in training deep models (Guo et al., 2018; Hacohen and Weinshall, 2019). In the NLP domain, Xu et al. (2020) first explore and validate the effectiveness of CL in the context of finetuning LMs on Natural Language Understanding (NLU) tasks. Their framework achieved general performance improvements across various NLU tasks, including Machine Reading Comprehension (MRC) and Natural Language Inference (NLI). In terms of retrievers, recent research has shown that by gradually increasing the difficulty of sampled data, CL can also bring significant performance improvements in the training of embedders (Zhu et al., 2022; Zeng et al., 2022; He et al., 2023).
3 Methodology
@@
In this section, we will provide a detailed introduction to our CL-RAG training framework. We first briefly introduce the RAG pipeline in Section 3.1. Then, in Sections 3.2 and 3.3, we present the curriculum learning methods we propose for RALM training and retriever training, respectively. An overview of our CL-RAG training framework is given in Figure 2.
3.1 Preliminary
@@
The RAG system combines a Retriever that retrieves query-relevant documents from external data bases with a Generator (LLM) that synthesizes responses from retrieved documents.
@@
$score_{i}=cos(E(q),E(d_{i})).$
Typically, we select the $top$-$k$ documents with the highest scores as the input to the generator.
@@
Given the $top$-$k$ retrieved documents, the goal of the Generator (LLM) is to utilize these external documents to better answer the question. Generally, the retrieved documents are concatenated with the query $q$ as contextual information, and then fed into the Generator to produce the answer:
$Output=LLM(d_{1}\oplus d_{2}\oplus...\oplus d_{k},q).$
In this paper, we employ a raw retriever to retrieve $n$ relevant documents. For generator (LLM) training, we construct curriculum-ordered data from easy to difficult examples by applying data augmentation and document rewriting to $top$-$k$ retrieved documents. Subsequently, for retriever training, we leverage the finetuned generator (RALM) to assess the quality of $n$ retrieved documents. We then perform reranking and sampling to construct curriculum-ordered training data to train the retriever to align with RALM’s preference.
3.2 Curriculum Learning for RALM
@@
In this section, we will present the data construction process for training the RALM, along with the corresponding curriculum learning framework.
@@
To construct data with increasing difficulty for curriculum learning, we first categorize and construct documents of various difficulty levels. Previous work (Yoran et al., 2023; Fang et al., 2024) has considered data augmentation and constructed documents with relevant noise, irrelevant noise, and counterfactual noise. However, they neglected to include simple documents that definitely contain the correct answers, which is essential during the initial training stage for the model to learn the fundamental ability to answer questions from documents that contain the answers.
@@
We first classify documents into three levels of difficulty: Easy, Common, and Hard. The Easy level contains at least two documents with correct answers to ensure the model can acquire preliminary answer extraction capabilities. The Common level follows the original retrieved documents, which may not contain the exact correct answers, thereby training the model’s ability to infer answers using documents of average quality. The Hard level includes the most challenging documents for the model, potentially containing irrelevant or even harmful noise, which trains the model’s ability to counteract disruptive information.
@@
Since the retriever’s results may not always include the correct answers, to construct the documents for Easy level, we use the documents provided for the corresponding questions in the MRQA reading comprehension task (Fisch et al., 2019) as the golden documents to ensure that they contain the exact correct answers. Meanwhile, we rewrite the golden document as shown in Figure 3 to obtain document that better meet the model’s needs. Compared to using only the golden document, using both the golden document and its rewritten version ensures that there are at least two documents containing answers within the $top$-$k$ documents, which helps train the model’s ability to extract consistent information from multiple documents. We complete the total number of documents to $k$ using the original retrieved documents.
@@
For Common level, we directly use the $top$-$k$ documents from the retriever’s results. Unlike documents in Easy level, these may not contain the correct answers and have more noise related to the questions, which may increase the difficulty for the model to answer the questions.
@@
For Hard level, we randomly select a document from the $top$-$k$ and replace it with retrieved document from other questions or perform counterfactual rewriting as shown in Figure 3. This introduces irrelevant and counterfactual noise, which represents the most challenging type of documents for the model, aiming to train the model’s robustness against various types of real-world data.
@@
3.3 Curriculum Learning for Retriever
In this section, we will introduce our retriever training strategy. It initially employs a well-trained RALM to rerank documents and constructs data through stratified sampling. Subsequently, the retriever is trained in stages using a curriculum learning approach to align with the RALM.
We first employ the trained RALM to assess documents, thereby eliciting its preferences for them. Differs from that in Lin et al. (2023), which relies on the raw LLM, utilizing RALM for evaluation provides a more effective means of differentiating between high-quality and low-quality documents.
Previous researchers (Shi et al., 2023; Zhang et al., 2024) assess document quality based on the probability of the model generating the correct answer and the improvement in decoding rank, respectively. However, using either method alone has its limitation: the probability of generating the correct answer can be unstable when the input documents are diverse, and the decoding rank cannot finely distinguish between good and bad documents when the model itself is capable of generating the correct answer, i.e., when most of the input documents have a decoding rank of 1 for the correct answer.

### v2

# DACL-RAG: Data Augmentation Strategy with Curriculum Learning for Retrieval-Augmented Generation

@@ Abstract
Retrieval-Augmented Generation (RAG) is an effective method to enhance the capabilities of large language models (LLMs). Existing methods typically optimize the retriever or the generator in a RAG system by directly using the top-$k$ retrieved documents. However, two key issues inherent in the training data constrain the effectiveness of this training paradigm: (1) across different queries, the top-$k$ retrieved documents vary greatly in content quality, with some providing valuable knowledge while others lack critical information or are even misleading, and training on such data in a purely random manner may impair the generator’s ability to extract key information; (2) for a given query, the limited set of $k$ documents often exhibits low discriminability, and training solely on them makes it difficult for the retriever to learn how to distinguish between relevant and irrelevant documents. To address these issues, we introduce DACL-RAG, a multi-stage RAG training framework that combines a multi-level Data Augmentation strategy with a multi-stage Curriculum Learning paradigm. The data augmentation strategy constructs comprehensive and diverse training sets with controllable difficulty levels through sample evolution, while the curriculum learning paradigm organizes them into progressive stages for training, ensuring stable and consistent improvements, thereby optimizing the overall performance and generalization of the RAG system more effectively. Our DACL-RAG framework demonstrates consistent effectiveness across four open-domain QA datasets, achieving performance gains of 2% to 4% over multiple advanced methods.
I Introduction
Figure 1: Existing training paradigm. Here, green denotes documents that support the model’s responses, while red denotes documents that are useless or even harmful. We conduct analysis on four open-domain QA datasets: NQ, TriviaQA, HotpotQA, and PopQA. For generator training (upper panel): Raw Retrieve Result reports the answer recall of the original retrieved documents; these raw retrievals exhibit low recall of answer-bearing evidence, with some documents containing the answer and others not, so training directly on such inputs injects noise and impedes optimization. For retriever training (lower panel): after scoring document usefulness with an LLM, we find that within the toptop-55 candidates the ranking scores exhibit limited dispersion, providing insufficient signal to learn robust discrimination between helpful and unhelpful documents; expanding the sampling window to the toptop-2020 increases score separation and mitigates this issue.
@@ I Introduction
Large language models (LLMs) have demonstrated remarkable capabilities in a wide range of Natural Language Processing (NLP) tasks [1, 2, 3], but they are still constrained by the limitations of the static knowledge embedded within their internal parameters [4, 5, 6]. Retrieval-Augmented Generation (RAG) addresses this limitation by supplementing LLMs with additional knowledge retrieved from external knowledge bases, and has significantly enhanced the capabilities of existing large models in tasks such as Open-Domain Question Answering [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17] and Dialog System [18, 19, 20].
The overall performance of the RAG system depends crucially on the quality of the retrieved documents and the LLMs’ ability to effectively utilize them. Accordingly, much effort has been devoted to enhancing RAG systems from two main directions: (1) improving the retriever by leveraging LLMs to assess document utility and provide supervision signals [11, 21, 22, 15, 16]; (2) training Retrieval-Augmented Language Models (RALMs) that can better utilize retrieved documents [7, 8, 9, 23, 14, 12, 24], or combining both [25, 13] to enhance the overall performance. These methods generally employ an off-the-shelf retriever to obtain the toptop-$k$ most relevant documents for each query, which are then combined with the corresponding query to train either the retriever or the generator separately. However, the effectiveness of this training paradigm is constrained by two key issues stemming from the inherent characteristics of the training data. (1) Across different queries, the toptop-$k$ retrieved documents vary greatly in content quality, which introduces considerable noise into generator training, as shown in Fig. 1 (top panel). Among the retrieved documents, the toptop-55 answer recall is 63.19%, meaning that many queries still have no answer in the top five retrieved documents, reflecting substantial noise and distractors. When such documents are randomly used during training, they degrade the learning process of the generator, hindering its ability to effectively extract and utilize relevant information. (2) For a given query, the toptop-$k$ documents often have comparable effectiveness in supporting the answer, thereby limiting informational diversity and weakening the supervision signal for retriever training. As shown in Fig. 1 (bottom panel), for over 50% of queries, the variance of answer-utility scores across the top-5 documents is less than 1, indicating limited dispersion and thereby making it difficult for the retriever to learn fine-grained distinctions in document contribution. A few recent studies have attempted to address these issues by synthesizing different types of hard-to-learn noise to improve the robustness of the generator [12, 14]. However, this approach neglect the importance of clean data, which is essential for enabling RALMs to extract and utilize relevant information effectively, and offer no benefit toward retriever optimization. Thus, how to construct a balanced and systematic training strategy to improve the learning of both retriever and generator remains underexplored and deserves more attention.
Curriculum learning [26] provides an effective data-organization scheme in which models begin with easy examples and gradually tackle more difficult ones, thereby building the capacity to cope with noise and ambiguous samples. Building on this principle, we introduce DACL-RAG, a novel RAG training framework that combines a multi-level Data Augmentation strategy with a multi-stage Curriculum Learning paradigm. The data augmentation strategy constructs comprehensive and diverse training sets with controllable difficulty levels, while the curriculum learning paradigm organizes them into progressive stages for training, ensuring stable and consistent improvements. Specifically, for the generator, difficulty is defined by the noise ratio within documents. By rewriting golden documents to enhance query information or craft counterfactual distractor documents, we construct multiple training datasets with gradually increasing noise ratios. For retriever training, difficulty is defined based on the discriminability in utility between different documents for answering a given query. The smaller the discriminability in utility, the higher the difficulty, as the limited separability provides insufficient signal for the retriever to learn to distinguish useful from unhelpful documents. Compared to existing methods that typically select only the top-ranked documents, we first broaden the overall sampling range to include a wider variety of documents. A well-trained RALM is used to evaluate utility of documents and re-rank them, and we construct training sets from easy to difficult by gradually reducing the ranking gap between sampled documents.
@@
As a result, DACL-RAG bring two advantages to RAG system: (1) For the generator, DACL-RAG systematically transitions it from merely extracting information to effectively countering potential distracting noise within documents. (2) For the retriever, DACL-RAG train it in stages to progressively distinguish documents with obvious differences and then distinguish those with only slight differences. This progressive training framework has led to an overall performance and generalization improvement of the RAG system. Our contributions can be concluded as follows:
@@
  • We propose the DACL-RAG training framework for RAG system based on the concept of curriculum learning. To the best of our knowledge, this is one of the first times that CL strategy has been applied to optimize RAG systems.
@@
  • We defined the task-specific difficulty levels for the retriever and generator in RAG, and designed a complete method for constructing curriculum-structured training sets.
@@
  • We evaluated our DACL-RAG framework on four popular datasets, demonstrating consistent performance gains of 2% to 4% over multiple advanced methods, thereby highlighting the superiority of our approach.
II Related Work
II-A Retrieval-augmented Generation
@@
Using documents retrieved from extended knowledge bases to enhance the capabilities of large language models (LLMs) has been proven effective in NLP tasks, including language modeling [27, 21] and question answering [7, 8, 9, 10, 11, 12, 13, 14, 15, 17]. Specifically, a Retrieval-Augmented Generator (RAG) system takes a query as input and uses a retriever to retrieve relevant documents from an external knowledge base. Then, it combines the documents with the query and feeds them into the LLM to make up for it’s own lack of knowledge.
Optimization of the RAG system focuses on two main areas: improving the retriever and enhancing the generator (LLM) to the RALM. On the retriever side, Replug [11] uses KL divergence to align retriever results with generator’s preferences. ARL2 [22] uses LLM to directly label evidence support and combines list-wise and pairwise ranking losses with adaptive self-training. LLM-Embedder [21] employs a distillation objective based on model rankings. SuRe [28] performs evidence summarization and assessment over multiple candidate answers to choose the model’s most endorsed output. RECOMP [16] trains a compressor to summarize documents returned by the retriever, thereby providing more suitable context for the generator. BGM [15] trains an intermediate model that aligns retrieved evidence with the generator’s preferences. In general, these approaches either optimize the retriever directly or employ an intermediate model to better match the generator’s preferences.
For the generator, RAG [7], AmbigQA [8] and BART [9] introduced retrieval augmented generative models for open domain question answering. FiD [23] finetunes the generator to handle retrieved documents and queries, addressing irrelevant information. LUMEN [29] adds an offline–online memory encoding scheme for FiD, whereas Glimmer [30] further adds a shallow reranker over the pre-computed memory. Self-RAG [31] trained the generator to choose whether and when to retrieve, leading to better answers. RankRAG [24] instruction-tunes a single generator to perform both context reranking and answer generation, improving retrieval–generation consistency. Other studies introduce noise to improve model robustness [12, 14].
Combining the strengths of both approaches, RA-DIT [13] uses modular training to optimize the retriever and generator separately, enhancing overall RAG system performance. RAG-DDR [25] uses an end-to-end loop of sampling, scoring, and training to alternately update the refinement and generation modules, align their preferences, and improve RAG performance.
II-B Curriculum Learning
Curriculum Learning (CL) [26] is a machine learning strategy that mimics human learning by training models gradually, starting with simpler tasks and progressing to more complex ones. It aims to improve the model’s generalization ability and accelerate its convergence speed [32]. Early studies have extensively investigated CL in computer vision domain and have demonstrated its advantages in training deep models [33, 34, 35, 36]. In the NLP domain, [37] adopt a heuristic approach to difficulty ordering, building the curriculum using sentence length and the rarity of vocabulary, however, empirical evidence shows that such heuristics yield only limited gains [38]. Beyond such heuristic annotation schemes, [39] systematically explore and validate Curriculum Learning (CL) for fine-tuning language models on Natural Language Understanding (NLU) tasks by leveraging external teacher models, achieving consistent gains across Machine Reading Comprehension (MRC), Natural Language Inference (NLI), and related benchmarks. Building on this line, [40] leverage the PLM itself to estimate sample difficulty and prioritize fine-tuning examples accordingly, improving both convergence speed and final performance. Recent work has further applied the CL paradigm to specialized domains such as code [41] and knowledge graphs [42]. For retrievers, recent work demonstrates that progressively increasing the difficulty of sampled data likewise yields substantial improvements when training embedding-based models [43, 44, 45].
Figure 2: The overview of our DACL-RAG training framework, which has into two continuous phases: (1) Training Generator: We construct multiple difficulty levels of documents and then finetune a RALM in a stage-by-stage manner. (2) Training Retriever: We use the well-trained RALM to assess documents and rerank them. We then construct document data from easy to difficult, and finetune the retriever in a stage-by-stage manner.
III Methodology
@@
In this section, we will provide a detailed introduction to our DACL-RAG training framework. We first briefly introduce the RAG pipeline in Section III-A. Then, in Sections III-B and III-C, we present the data augmentation and curriculum learning methods we propose for RALM training and retriever training, respectively. An overview of our DACL-RAG training framework is given in Figure 2.
III-A Preliminary
@@
The RAG system combines a Retriever that retrieves query-relevant documents from external data bases with a Generator that synthesizes responses from retrieved documents.
@@
$score_{i}=cos(E(q),E(d_{i})).$ (1)
Typically, we select the toptop-$k$ documents with the highest scores as the input to the generator.
@@
Given the toptop-$k$ retrieved documents, the goal of the Generator (i.e., the LLM) is to utilize these external documents to better answer the question. Generally, the retrieved documents are concatenated with the query $q$ as contextual information, and then fed into the Generator to produce the answer:
$Output=LLM(d_{1}\oplus d_{2}\oplus...\oplus d_{k},q).$ (2)
In this paper, we first employ a raw retriever to retrieve $n\,(n\gg k)$ relevant documents for each query. For generator (LLM) training, we construct curriculum-ordered data from easy to difficult examples by applying data evolution and document rewriting to toptop-$k$ retrieved documents. For retriever training, we leverage the well-trained generator (RALM) to assess the quality of $n$ retrieved documents. We then perform reranking and sampling to construct curriculum-ordered training data to train the retriever to align with RALM’s preference.
III-B Data Augmentation and Curriculum Learning for RALM
@@
To improve the model’s robustness in answering queries with noisy documents, we define the difficulty level of training sets according to the ratio of noise in the documents. Previous work [12, 14] has considered data augmentation and constructed documents with relevant noise, irrelevant noise, and counterfactual noise. However, they neglected to include simple documents that definitely contain the correct answers, which is essential during the initial training stage for the model to learn the fundamental ability to answer questions from documents that contain the answers.
@@
We first define three levels of difficulty for documents: Easy, Common, and Hard. The Easy level contains at least two documents with correct answers to ensure the model can acquire preliminary answer extraction capabilities. The Common level follows the original retrieved documents, which may not contain the exact correct answers, thereby training the model’s ability to infer answers using documents of average quality. The Hard level includes the most challenging documents for the model, potentially containing irrelevant or even harmful noise, which trains the model’s ability to counteract disruptive information.
@@
Since the retriever’s results may not always include the correct answers, to construct the documents for Easy level, we use the documents provided for the corresponding questions in the MRQA reading comprehension task [46] as the golden documents to ensure that they contain the exact correct answers. Meanwhile, we rewrite the golden document as shown in the upper portion of Figure 3 to obtain document that better meet the model’s needs. Our rewriting approach preserves all key information necessary for answering questions while eliminating redundant segments from the original golden documents. Compared to using only the golden document, using both the golden document and its rewritten version ensures that there are at least two documents containing answers within the toptop-$k$ documents, which helps train the model’s ability to extract consistent information from multiple documents. The improvements brought by this approach are empirically validated through ablation studies in section V-C. We then complete the total number of documents to $k$ using the original retrieved documents.
@@
For Common level, we directly use the toptop-$k$ documents from the retriever’s results. Unlike documents in Easy level, these may not contain the correct answers and have more noise related to the questions, which may increase the difficulty for the model to answer the questions.
@@
For Hard level, we randomly select a document from the toptop-$k$ and replace it with retrieved document from other questions or perform counterfactual rewriting as shown in the lower portion of Figure 3. This introduces irrelevant and counterfactual noise, which represents the most challenging type of documents for the model, aiming to train the model’s robustness against various types of real-world data.
To demonstrate the difficulty differences across different levels, we test each level of data using a naive RAG system (with Contriever-MSMARCO as the retriever and LLaMA3-8B-Instruct as the generator). The results are shown in Table I. We established the document difficulty assessment framework based on two perspectives and four key metrics: (1) For LLM generation performance, we evaluate using Exact Match (EM) and F1 scores. (2) For inherent document quality, we evaluate using the recall rate (R@k) of documents (here we use both R@1 and R@5). The definitions of these metrics are detailed in Section IV-B. The difficulty assessment results are shown in Table I, where we observe significant differences in difficulty levels across various training stages.
@@
III-C Data Augmentation and Curriculum Learning for Retriever
TABLE I: Evaluation of Document Data Difficulty Across Different Generator Training Stages
Difficulty Stages EM F1 R1 R5
Easy 1 61.23 74.33 65.38 100.00
Common 2 47.87 59.77 38.00 59.78
Hard 3 41.67 52.99 20.17 54.68
In this section, we will introduce our retriever training strategy. It initially employs a well-trained RALM to rerank retrieved documents and, via stratified sampling, construct three difficulty levels: easy, common, and hard, where simple corresponds to larger inter-document quality gaps for the retriever, and hard corresponds to smaller gaps. The retriever is then trained in stages, following a curriculum learning regime that progressively increases difficulty and aligns the retriever with the RALM.
We first employ the trained RALM to assess documents, thereby eliciting its preferences for them. Differs from that in [13], which relies on the raw LLM, utilizing a well-trained RALM for evaluation provides a more effective means of differentiating between high-quality and low-quality documents.
Previous researchers [11, 21] assess document quality based on the probability of the model generating the correct answer and the improvement in decoding rank, respectively. However, using either method alone has its limitation: the probability of generating the correct answer can be unstable when the input documents are diverse, i.e., there may exist multiple documents containing answers, but due to their diverse content, the probability of generating accurate answers varies significantly, making it difficult to reliably assess document quality. On the other hand, relying solely on the decoding rank cannot finely distinguish between good and bad documents when the model itself is capable of generating the correct answer, i.e., when most of the input documents have a decoding rank of 1 for the correct answer, indicating gains from the model’s inherent capability rather than document quality.